You have taken a long position in a call option on IBM common stock. The option has an exercise price of $147 and IBM's stock currently trades at $152. The option premium is $6 per contract. a. How much of the option premium is due to intrinsic value versus time value? Option Premium Intrinsic value $ Time value b. What is your net profit on the option if IBM’s stock price increases to $162 at expiration of the option and you exercise the option? (Negative amount should be indicated by a minus sign.)
Answer #1

Given for IBM stock,

Strike price of call option = $147

spot price = $152

Option premium = $6

Intrinsic value of a call option = Spot price - strike price = 152-147 = $5

time value of a call option = option premium - intrinsic value = 6-5 = $1

If at expiration, stock price = $162

total profit = stock price - strike price - option premium = 162 - 147 - 6 = $9 per option.